OpenStudy (anonymous):

Choose the graph which matches the following system of equations. 5x + 2y = 2 3x – 3y = 18

OpenStudy (nick1232):

What graph

OpenStudy (anonymous):

hold i'm posting it

OpenStudy (nick1232):

alright

OpenStudy (anonymous):

@Nick1232 can u help me?

OpenStudy (nick1232):

5x + 2y = 2 2y=5x+2 /2 /2 /2 y=-5/2+1

OpenStudy (nick1232):

Then graph that

OpenStudy (anonymous):

The graph similar to that one.

OpenStudy (nick1232):

3x – 3y = 18 -3y= 3x+2 /-3 /-3 /-3 y= -1 - .67

OpenStudy (nick1232):

Then graph that

OpenStudy (anonymous):

is it b?

OpenStudy (nick1232):

Graph both of them and you will find your answer

OpenStudy (anonymous):

is it b?

OpenStudy (anonymous):

@Nick1232 is it b?

OpenStudy (anonymous):

Yup.

OpenStudy (nick1232):

idk i have to graph it

OpenStudy (nick1232):

But im on my computer

OpenStudy (nick1232):

lemme see what i can do

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

If it intercept at x = 2, y = -4 then it's the answer.

OpenStudy (nick1232):

hold up i almost have it , its not A for sure

OpenStudy (anonymous):

@DontLikeMathButOhWell thanks for the help

OpenStudy (anonymous):

@Nick1232 okay

OpenStudy (nick1232):

or B !

OpenStudy (anonymous):

its not b?

OpenStudy (nick1232):

Dang its so hard to do it from a computer , im going to write it real quick ok ?

OpenStudy (anonymous):

@Nick1232 okay

OpenStudy (anonymous):

@Nick1232 did u get it yet?

OpenStudy (nick1232):

D

OpenStudy (anonymous):

thanks a lot =D @Nick1232

OpenStudy (nick1232):

no not d , imma look at your graph again

OpenStudy (anonymous):

but r u sure?

OpenStudy (anonymous):

ohhh ok

OpenStudy (nick1232):

Need to check my work

OpenStudy (anonymous):

ok

OpenStudy (nick1232):

Alright its C im positive

OpenStudy (anonymous):

u sure?

OpenStudy (nick1232):

100

OpenStudy (anonymous):

okay i have 2 more problems left and i'll let u know if its correct or not but thanks for ur help =)

OpenStudy (nick1232):

Alright , this is algebra 1 right?

OpenStudy (nick1232):

I just got out of there and im in geometry D:

OpenStudy (anonymous):

no its algebra 2

OpenStudy (nick1232):

Ooh seems like algebra 1

OpenStudy (nick1232):

I need help with Geometry

OpenStudy (anonymous):

ohh i only have algebra 2

OpenStudy (nick1232):

YOu never did geometry ?

OpenStudy (anonymous):

no i did but its just that i'm lucky i passed with a D cuz i really suck at math

OpenStudy (nick1232):

ahh dang alright

OpenStudy (anonymous):

Nick how did you get C?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

I worked out my graph and it became identical to choice B..I even put it in a solver and got the same looking graph..

OpenStudy (nick1232):

Thats weird , i did it my self and got C

OpenStudy (anonymous):

I even found another thread with the same question.. http://openstudy.com/study#/updates/4f09e669e4b014c09e643c66 but oh well if she got the answer right as C it's okay then.